The distance between the dipole and the point char
Answer #1

field of a dipole varies as 1/d3 on the dipole axis. whn d increase to 10d force reduces by 103 =1000.

Force between q and +Q , \propto \frac{1}{(d+s/2)^{2}}

Force between -q and +Q is, \propto \frac{1}{(d-s/2)^{2}}

netforce is the difference,\propto \frac{ds}{(d^{2}-s^{2}/4)^{2}}

\propto \frac{d}{(d^{2})^{2}} d>>s

\propto \frac{1}{d^{3}}